9. Transfer payments are: A. spending that transfers resources from the government to individuals. B. payments that individuals make to the government. C. when individuals transfer stock ownership in lieu of payment with cash or other liquid assets. D. when individuals transfer cash for payments of a good or service.

10. U.S. exports are: A. U.S. goods sold to foreigners. B. Foreign goods bought by Americans. C. U.S. goods sold to Americans. D. Foreign and U.S. goods sold to foreigners, but consumed in the U.S.

11. Suppose that in 2018 a country imported $100M worth of goods and services and exported $125M worth of goods and services. In this case, this country had a _____ of _____. A. trade deficit; $25M B. trade deficit; $100M C. trade surplus; $25M D. trade surplus; $125M 12. If consumption is $8 billion, investments is $4 billion, government purchases are $2 billion, imports are $1 billion, and exports are $2 billion, GDP must equal: A. $16 billion. B. $17 billion. C. $15 billion. D. $14 billion.

Consider an economy with the following market prices and production quantities over time: Good A Good B Good C Year P ($) Q (units) P ($) Q (units) P ($) Q (units) 2017 9.00 10 5.00 56 21.00 40 2018 9.00 12 5.25 60 22.00 42

13. Nominal GDP in 2017 was _____. A. $1,210 B. $1,347 C. $1,425 D. None of the above/not enough information

14. Nominal GDP in 2018 was _____. A. $1,180 B. $1,347 C. $1,425 D. None of the above/not enough information

15. Using 2018 as the base year, real GDP in 2017 was _____. A. $1,264 B. $1,347 C. $1,425 D. None of the above/not enough information

16. Using 2018 as the base year, real GDP in 2018 was _____. A. $1,290 B. $1,347 C. $1,425 D. None of the above/not enough information
Answer #1

9. A. spending that transfers resources from the government to individuals.
(Transfer payments are transfer of resources from government to individuals.)

10. A. U.S. goods sold to foreigners.
(Exports of a country are goods of that country sold to foreigners)

11. C. trade surplus; $25M
(Trade balance = Exports - Imports = 125-100 = 25 so there is a surplus as trade balance is positive.)

12. C. $15 billion.
(GDP = C + I + G + X - M = 8 + 4 + 2 + 2 - 1 = 15)

(Post 4 MCQs at a time.)